Madylencatgirl:

The volume of a sphere whose diameter is 18 centimeters is how many π cubic centimeters. If its diameter were reduced by half, its volume would be of its original volume.

563blackghost:

Volume of a sphere's formula is: \(\large\bf{V=\frac{4}{3} \pi r^{3}}\) You need to know the radius not diameter. You can divide the diameter by 2 to get the radius. \(\bf{18 \div 2 = 9}\) You then plug this into its formula. `REMEMBER TO KEEP PI AS π` \(\large\bf{V=\frac{4}{3} \pi 9^{3}}\) Simplify.